# FirstReferences

### v2

@@ Appendix G VAET-CT crossover
At small $g/\omega=0.1$, the system undergoes VAET dynamics, showing the expected sharp transfer resonance at $\Delta E=\sqrt{\omega^{2}-(2V)^{2}}$ from the single-phonon assisted process. As $g/\omega$ increases, the vibrational state of the system becomes effectively displaced, giving rise to non-negligible transfer rates at resonances associated with higher-order phonon-assisted processes. However, when $g\sim\omega$, the displaced LVCM system becomes an adiabatic CT system, leading to monotonically increasing transfer rates at low $\Delta E$ and sharp resonances at higher $\Delta E$ (see the results with $g/\omega=0.5$ and 1 in Fig. S6B). From these observations, we suppose that $g\approx 0.5\omega$ is the crossover between VAET and CT. Therefore, in this work, we associate systems with $g_{i}/\omega_{i}\lesssim 0.5$ to non-perturbative VAET and systems with $g_{i}/\omega_{i}\gtrsim 1$ to non-perturbative CT, all with strong electronic coupling ($|V|\sim\lambda_{i}/4$).
As we further increase $g/\omega$, the system progressively evolves into the nonadiabatic regime of CT, where discernible resonance peaks are recovered across the transfer rate spectrum. Differently from VAET, these resonances occur at integer multiples of $\omega$, and the position of the strongest resonance depends on the displacement between the donor and acceptor potential energy surfaces. The two distinct regimes within CT, nonadiabatic and adiabatic transfers, are determined by the relative strengths of the electronic coupling to the reorganization energy of the system and the dissipation rate of the vibrational mode, which have been studied in Refs. [28, 20].

### v3

@@ Appendix G VAET-CT crossover
Figure S7: CT rates for single-mode systems. (A) Transfer rate spectrum with $(V,g_{1},\gamma_{1})=(0.200,1.200,0.036)\omega$. The solid blue curve shows the transfer rate calculated from Eq. (4), using the definition in Eq. (5) and including spin decoherence ($\gamma_{z}\!=\!0.0014\omega$) and motional dephasing ($\gamma_{1m}\!=\!0.0160\omega$). (B) Numerical donor population evolution $P_{D}(t)$ versus energy gap $\Delta E$ and the number of vibrational oscillations $\omega t/2\pi$, using the same parameters as the solid red curve in (A). (C and D) Same layout as in (A and B) with $(V,g_{1},\gamma_{1},\gamma_{z},\gamma_{1m})=(0.138,1.029,0.023,0.0009,0.0100)\omega_{1}$.
At small $g/\omega=0.1$, the system undergoes VAET dynamics, showing the expected sharp transfer resonance at $\Delta E=\sqrt{\omega^{2}-(2V)^{2}}$ from the single-phonon assisted process. As $g/\omega$ increases, the vibrational state of the system becomes effectively displaced, giving rise to non-negligible transfer rates at resonances associated with higher-order phonon-assisted processes. However, when $g\sim\omega$, the displaced LVCM system becomes a CT system, particularly an adiabatic CT system with $|V|\sim\lambda/4$, leading to monotonically increasing transfer rates at low $\Delta E$ and sharp resonances at higher $\Delta E$ (see the results with $g/\omega=0.5$ and 1 in Fig. S6B). From these observations, we infer that $g\approx 0.5\omega$ marks the crossover between VAET and CT, where the first resonant peak associated with a single-phonon assisted process is effectively washed out into a monotonically increasing feature in the transfer rate. Therefore, in this work, we associate systems with $g_{i}/\omega_{i}\lesssim 0.5$ to non-perturbative VAET and systems with $g_{i}/\omega_{i}\gtrsim 1$ to non-perturbative CT, all with strong electronic coupling ($|V|\sim\lambda_{i}/4$).
As we further increase $g/\omega$, the system progressively evolves into the nonadiabatic regime of CT with $|V|<\lambda/4$, where discernible resonance peaks are recovered across the transfer rate spectrum. Different from VAET, these resonances occur at integer multiples of $\omega$, and the position of the strongest resonance depends on the displacement between the donor and acceptor potential energy surfaces. The two distinct regimes within CT, nonadiabatic and adiabatic transfers, are determined by the relative strengths of the electronic coupling to the reorganization energy of the system and the dissipation rate of the vibrational mode, which have been studied in Refs. [26, 20].
Appendix H Single-mode charge transfer dynamics
Fig. S7 shows the transfer rate spectra and the corresponding donor population dynamics from numerical simulations for the single-mode cases, which are used for comparison with the two-mode CT results in Fig. 3 of the main text.